A person plans to invest a total of ​$190,000 in a money market​ account, a bond​ fund, an international stock​ fund, and a domestic stock fund. She wants 60​% of her investment to be conservative​ (money market and​ bonds). She wants the amount in domestic stocks to be 44 times the amount in international stocks.​ Finally, she needs an annual return of ​$7,600. Assuming she gets annual returns of​ 2.5% on the money market​ account, 3.5% on the bond​ fund, 4% on the international stock​ fund, and​ 6% on the domestic stock​ fund, how much should she put in each​ investment?
Answer by addingup(3677)

Total to invest: 190,000
60% in Money Market and Bond Funds:
190,000*0.6 = 114,000 in MM and BF
~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~
Interest:
MM 0.025
BF 0.035
IS 0.04
DS 0.06
~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~
MM+BF = 114,000
IS+DS = 190,000-114,000 = 76,000
~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~
DS = 44*IS
DS+IS = 76,000
44*IS+IS = 76,000; 45IS = 76,000; IS = 76,000/45 = 1,688.89
IS = 1,688.89
DS = 44*1,688.89 = 74,311.16
~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~
IS = 1,688.89*0.04 = $67.55  interest on International Stocks
DS = 74,311.16*0.06 = $4,458.67  interest on Domestic Stocks
Total: 67.55+4,458.67 = 4,526.22
~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~
MM+BF = 114,000
0.025MM+0.035BF = 7,600-4,526.22 = 3,073.78
0.025x+0.035(114,000-x) = 3,073.78 = 91,622  Invested at 2.5%
114,000-91,622 = 22,378  Invested at 3.5%
~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~
91,622*0.025 = 2,290.55 Interest from Money Market
22,378*0.035 = 783.23 Interest from Bond Funds
~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~ ~
Check:
67.55+4458.67+2290.55+783.23 = 7,600 Correct.
:
John